# Decay and production properties of strange double charm pentaquark

### v2

@@ Abstract
In this work we investigate the decay and production properties of the strange double-charm pentaquark $P_{ccs}^{++}$ with strangeness S=1S=-1. Building upon our previous work predicting its $J^{P}=1/2^{-}$ molecular configuration, we employ three-point QCD sum rules to calculate its strong decay widths and estimate its production branching ratios via $\Xi_{bc}^{+}$ baryon decays. The total strong decay width into the $\Xi_{cc}\bar{K}$ and $\Omega_{cc}\pi$ final-state channels is determined as $84.58^{+19.25}_{-18.80}$ MeV. Furthermore, using a rescattering mechanism, we analyze the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process and estimate the production branching ratio to be $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})=(4.32_{-1.47}^{+2.02})\times 10^{-6}$. The relatively narrow width and detectable branching ratio suggest that this pentaquark state could be observed in experiments such as LHCb.
@@ I Introduction
The study of exotic multiquark states, proposed early in 1964 Gell-Mann:1964ewy ; 1964-Zweig-p- , has become a pivotal frontier in hadronic physics, offering profound insights into the nonperturbative dynamics of quantum chromodynamics (QCD) Nielsen:2009uh ; Chen:2016qju ; Richard:2016eis ; Esposito:2016noz ; Ali:2017jda ; Guo:2017jvc ; Albuquerque:2018jkn ; Liu:2019zoy ; Brambilla:2019esw ; Richard:2019cmi ; Faustov:2021hjs ; Chen:2022asf ; Meng:2022ozq . Since the discovery of the first hidden-charm pentaquarks $P_{c}(4380)$ and $P_{c}(4450)$ by the LHCb Collaboration LHCb:2015yax , significant theoretical and experimental efforts have been devoted to unraveling the nature of these states, which lie beyond the conventional quark model. The recent observation of the double-charm tetraquark $T_{cc}^{+}(3875)$ LHCb:2021vvq ; LHCb:2021auc and the strange-charm tetraquark $T_{c\bar{s}}(2900)$ LHCb:2022lzp further highlights the rich spectrum of exotic hadrons and underscores the potential existence of their double-heavy counterparts, such as double-charm pentaquarks.
@@
Meanwhile, the ongoing experimental quest for doubly heavy $\Xi_{bc}^{+}$ baryons offers a pivotal opportunity to unravel the double-charm pentaquark state. Cabibbo-favored decays of $\Xi_{bc}^{+}$, such as $\Xi_{bc}^{+}\to D_{s}^{*-}\Xi_{cc}^{++}$, could generate the $\Xi_{cc}\bar{K}$ pentaquark through rescattering processes. At the quark level, the diagram for the process $\Xi_{bc}^{+}\rightarrow D^{-}P^{++}_{ccs}$ is shown in the left panel of Fig. 1. The weak decay arises from a Cabibbo-favored weak transition $b\rightarrow c(\bar{c}s)$ along with the creation of a $d\bar{d}$ pair from the strong interaction. This diagram, known as the external W emission diagram, is non-factorizable because the $s\bar{c}$ pair produced in the weak interaction ends up in different final-state hadrons. Thus, long-distance contributions play a significant role in the $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ process, where the weakly produced $c\bar{s}$ and uccucc pairs hadronize as $D_{s}^{\ast-}$ and $\Xi_{cc}^{++}$, respectively, followed by a strong rescattering between $D^{-}$ and $P_{ccs}^{++}$ mediated by $\bar{K}^{0}$ meson exchange. The corresponding rescattering process at the hadronic level is shown in the right panel of Fig. 1. The rescattering mechanism for final-state-interaction (FSI) effects has been successfully applied to $D$ meson decays Li:2002pj , $B$ meson decays Cheng:2004ru ; Lu:2005mx , and charm baryon decays Han:2021azw ; Jia:2024pyb . Recently, this mechanism has also been applied to the production of tetraquark states in $B$ meson decays Chen:2020eyu ; Yang:2024coj and of hidden-charm pentaquark states $P_{c}$ in $\Xi_{b}$ decays Hsiao:2024szt . In this work, we apply the rescattering mechanism to the production of the double-charm pentaquark via the $\Xi_{bc}^{+}\rightarrow D^{\ast-}_{s}\Xi_{cc}^{++}\rightarrow D^{-}P^{++}_{ccs}$ process with $\bar{K}^{0}$ exchange.
@@
A precise calculation of the decay width and production branching ratios of the double-charm pentaquark is thus critical to quantify its production rates in such channels and to guide experimental searches at facilities like LHCb and Belle II. In this work, we continue our study by calculating the decay width of the double-charm pentaquark using the QCD sum rule method and estimating its production branching ratios in $\Xi_{bc}$ decays. This paper is organized as follows: In Sec. II, we outline the formalism for computing decay widths within the QCD sum rule approach. Sec. III presents the numerical results for the $J^{P}=1/2^{-}$ pentaquark’s decay properties. Sec. IV discusses its production mechanism via $\Xi_{bc}^{+}$ decays and estimates detectable branching ratios. A brief summary is presented in Sec. V.
@@ II Three Point QCD Sum Rule
Over past several decades, the method of QCD sum rule has been proven to be very powerful to study hadron properties Reinders:1984sr ; Shifman:1978bx ; Colangelo:2000dp ; Narison:2002woh . In this section, we shall study the three-point correlation function of several two-body strong decay process $M\rightarrow X+Y$. For the strong decay process $M\rightarrow X+Y$, the corresponding correlator is written as
@@
where $J_{M(X,Y)}$ is the interpolating current for the initial(final) state. In this section, we shall consider the $P_{ccs}\Xi_{cc}\bar{K}$ and $P_{ccs}\Omega_{cc}\pi$ strong decay vertices with $K(\pi)$ off shell. We use the following interpolating currents for $P_{ccs}^{++}$ by considering it as a $\Xi_{cc}\bar{K}$ molecule Yang:2024okq :
@@
where $\mathcal{C}$ denote the charge conjugate operator, subscript $a\cdots d$ denote the color index and $u,d,s,c$ denote the up, down, strange, charm quark field, respectively. These current can couple to the $P_{ccs}^{++}$ state with $J^{P}=1/2^{-}$ via
@@
which can coupling to the meson states via
@@
which can coupling to the baryon states via
@@
On the OPE side, we can evaluate the correlation function with standard QCD sum rule approach. To establish a sum rule for the coupling constant, we will pick out the $1/q^{2}$ terms around the pole $q^{2}\sim 0$ with the structure $\not${p} in the OPE series and then match both sides of the sum rule. To apply sum rules appropriately, we shall calculated at $Q^{2}$ far away from the on-shell mass $m_{K}^{2}$ to ensure the approximation $p^{2}=p^{{}^{\prime}2}=P^{2}$ valid. After performing the Borel transform $P^{2}\rightarrow M_{B}^{2}$ on both phenomenological and OPE sides, we obtain the strong coupling for $P_{ccs}\Xi_{cc}\bar{K}$ vertex
$g_{P_{ccs}\Xi_{cc}\bar{K}}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Xi_{cc}}\lambda_{K}(m_{P_{ccs}}+m_{\Xi_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Xi_{cc}}^{2}}{\mathrm{e}^{-m_{\Xi_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\frac{Q^{2}+m_{K}^{2}}{Q^{2}}\right)\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (13)
@@
$g_{P_{ccs}\Omega_{cc}\pi}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Omega_{cc}}\lambda_{\pi}(m_{P_{ccs}}+m_{\Omega_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Omega_{cc}}^{2}}{\mathrm{e}^{-m_{\Omega_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\frac{Q^{2}+m_{\pi}^{2}}{Q^{2}}\right)\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (14)
@@
Using the operator product expansion (OPE) method, the three-point function can also be evaluated at the quark-gluonic level as a function of various QCD parameters. To evaluate the Wilson coefficients, we adopt the quark propagator in momentum space and the propagator
@@
where $Q$ represents the heavy quark $c$ or $b$, $q$ represents the light quark $u,d,s$, the superscripts $a,b$ denote the color indices. In this work, we will evaluate Wilson coefficients of the correlation function up to dimension nine condensates at the leading order in $\alpha_{s}$. The spectrum function $\rho(s)$ in Eqs. (13), (14) are given in Appendix A. We shall discuss the detail to obtain suitable parameter working regions in QCD sum rule analysis in next section.
@@ III Numerical Analysis
where $t=\mathrm{log}\frac{\mu^{2}}{\Lambda^{2}}$, $b_{0}=\frac{33-2n_{f}}{12\pi}$, $b_{1}=\frac{153-19n_{f}}{24\pi^{2}}$, $b_{2}=\frac{2857-\frac{5033}{9}n_{f}+\frac{325}{27}n_{f}^{2}}{128\pi^{3}}$, $\Lambda=$210 MeV, 292 MeV and 332 MeV for the flavors $n_{f}=$5, 4 and 3, respectively. In this work, we evolve all the input parameters to the energy scale $\mu=m_{c}$ for our sum rule analysis. The parameters for the $K$ and $\pi$ mesons and double charm baryons are adopted in Tab. 1.
@@ III.1 Strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$
Figure 2: The dependence of the strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$ on the Borel mass $M_{B}^{2}$ (left panel) and transfer momentum $Q^{2}$ (right panel). On the left panel, the transfer momentum is set to be $Q^{2}=m_{\Xi_{cc}}^{2}\sim 13.1\;\mathrm{GeV}^{2}$. On the right panel, the red dots denote the value from Eq. (13) with $s_{0}=22.3\;\mathrm{GeV}^{2}$ and $M_{B}^{2}=1.77\;\mathrm{GeV}^{2}$. The blue solid line is the exponential fitting curve. The two dashed blue lines denote the upper and lower boundary of the uncertainty from various condensates, quark masses and hadronic parameters.
In the left panel of Fig. 2, we show the variation of the coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Xi_{cc}}^{2}\sim 13.1\;\mathrm{GeV}^{2}$. Such a momentum point is chosen far away from $m_{K}^{2}$ so that it can be safely ignored and the OPE series is valid in this region. We find that the coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$ has a maximum value at $M_{B}^{2}\sim 1.77\;\mathrm{GeV}^{2}$, around which it has minimal dependence on the non-physical parameter $M_{B}^{2}$. To extrapolate the coupling constant from the valid QCD sum rule working region to the physical pole $Q^{2}=-m_{K}^{2}$, we fit the sum rule result for $s_{0}=22.3\;\mathrm{GeV}^{2}$ and $M_{B}^{2}=1.77\;\mathrm{GeV}^{2}$ with exponential model
$g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})=a\,\mathrm{e}^{-b\,Q^{2}}.$ (18)
The fitting curve is shown in the right panel of Fig. 2, and the result is as follow:
$g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})=-(0.449\pm 0.054\;\mathrm{GeV}^{-3})\mathrm{e}^{-(0.001\pm 0.000)\;\mathrm{GeV}^{-2}\;Q^{2}}.$ (19)
With the on-shell condition $Q^{2}=-m_{K}^{2}$, we can obtain the strong coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(-m_{K}^{2})=-(0.45\pm 0.05)\;\mathrm{GeV}^{2}$. From the matrix element, we can obtain the decay width for $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Xi_{cc}}^{2},m_{K}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Xi_{cc}\bar{K}}(-m_{K}^{2})\left((m_{P_{ccs}}+m_{\Xi_{cc}})^{2}-m_{K}^{2}\right).$ (20)
Substitute the above on-shell coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=65.02\pm 15.69\,\mathrm{MeV}.$ (21)

### v4

@@ Abstract
In this work we investigate the decay and production properties of the strange double-charm pentaquark $P_{ccs}^{++}$ with strangeness S=1S=-1. Building upon our previous work predicting its $J^{P}=1/2^{-}$ molecular configuration, we employ three-point QCD sum rules to calculate its strong decay widths and estimate its production branching ratio via $\Xi_{bc}^{+}$ baryon decays. The total strong decay width to the $\Xi_{cc}\bar{K}$ and $\Omega_{cc}\pi$ final-state channels is determined as $85\pm 19$ MeV. Furthermore, using a rescattering mechanism, we analyze the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process and estimate the production branching ratio to be $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})=(4.3_{-1.5}^{+2.0})\times 10^{-6}$. The relatively narrow width and detectable branching ratio suggest the possibility searching for this pentaquark state in the future.
@@ I Introduction
The study of exotic multiquark states, proposed early in 1964 Gell-Mann:1964ewy ; 1964-Zweig-p- , has become a pivotal frontier in hadronic physics, offering profound insights into the nonperturbative dynamics of quantum chromodynamics (QCD) Nielsen:2009uh ; Chen:2016qju ; Richard:2016eis ; Esposito:2016noz ; Ali:2017jda ; Guo:2017jvc ; Albuquerque:2018jkn ; Liu:2019zoy ; Brambilla:2019esw ; Richard:2019cmi ; Faustov:2021hjs ; Chen:2022asf ; Meng:2022ozq . Since the discovery of the first hidden-charm pentaquarks $P_{c}(4380)$ and $P_{c}(4450)$ by the LHCb Collaboration LHCb:2015yax , significant theoretical and experimental efforts have been devoted to unraveling the nature of these states, which lie beyond the conventional quark model. The recent observation of the double-charm tetraquark $T_{cc}^{+}(3875)$ LHCb:2021vvq ; LHCb:2021auc and the strange-charm tetraquark $T_{c\bar{s}}(2900)$ LHCb:2022lzp further highlights the rich spectrum of exotic hadrons and underscores the potential existence of double-heavy counterparts in the baryon sector, such as double-charm pentaquarks.
@@
Meanwhile, the ongoing experimental quest for doubly heavy $\Xi_{bc}^{+}$ baryons offers a pivotal opportunity to unravel the double-charm pentaquark state. Cabibbo-favored decays of $\Xi_{bc}^{+}$, such as $\Xi_{bc}^{+}\to D_{s}^{*-}\Xi_{cc}^{++}$, could generate the $\Xi_{cc}\bar{K}$ pentaquark through rescattering processes. At the quark level, the diagram for the process $\Xi_{bc}^{+}\rightarrow D^{-}P^{++}_{ccs}$ is shown in the left panel of Fig. 1. The weak decay arises from a Cabibbo-favored weak transition $b\rightarrow c(\bar{c}s)$ along with the creation of a $d\bar{d}$ pair from the strong interaction. This diagram, known as the external W emission diagram, is non-factorizable because the $s\bar{c}$ pair produced in the weak interaction ends up in different final-state hadrons. Thus, long-distance contributions play a significant role in the $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ process, where the weakly produced $c\bar{s}$ and uccucc pairs hadronize via $\bar{K}^{0}$ meson exchange interaction, the $D_{s}^{\ast-}$ and $\Xi_{cc}^{++}$ pair converts into a $D^{-}$ and the pentaquark $P_{ccs}^{++}$. The corresponding interaction at the hadronic level is shown in the right panel of Fig. 1. Such a mechanism for final-state-interaction (FSI) effects has been successfully applied to $D$ meson decays Li:2002pj , $B$ meson decays Cheng:2004ru ; Lu:2005mx , and charm baryon decays Han:2021azw ; Jia:2024pyb . Recently, this mechanism has also been applied to the production of tetraquark states in $B$ meson decays Chen:2020eyu ; Yang:2024coj and of hidden-charm pentaquark states $P_{c}$ in $\Xi_{b}$ decays Hsiao:2024szt . In this work, we apply this mechanism to the production of the double-charm pentaquark via the $\Xi_{bc}^{+}\rightarrow D^{\ast-}_{s}\Xi_{cc}^{++}\rightarrow D^{-}P^{++}_{ccs}$ process with $\bar{K}^{0}$ exchange.
@@
A precise calculation of the decay width and production branching ratios of the double-charm pentaquark is thus critical to quantify its production rates in such channels and to guide experimental searches at facilities like LHCb and Belle II. In Ref. Yang:2024okq , we also predicted the masses of possible compact pentaquark states with the same quark content. Notably, two compact configurations were found with masses of $4.25\pm 0.4\;\text{GeV}$, which are close to the mass of the molecular state studied here. It is therefore a natural question how their decay widths and production rates compare with those of the molecular configuration. However, our calculations for these particular compact interpolating currents show that the sum rules analyses for the coupling constants $g_{P_{ccs}\Xi_{cc}\bar{K}}$ and $g_{P_{ccs}\Omega_{cc}\pi}$ did not exhibit a sufficiently stable Borel platform within a reasonable $M_{B}^{2}$ window, leading unreliable predictions of these couplings. Since the hadron decay widths and production rates are highly sensitive to the hadronic couplings, it is difficult to provide quantitative estimations for these compact pentaquark states within the framework of three-point QCD sum rules. In this work, we focus on the well-behaved molecular configuration to study the decay and production properties of strange double charm pentaquark state. We continue our study by calculating the decay width of the double-charm pentaquark using the QCD sum rule method and estimating its production branching ratios in $\Xi_{bc}$ decays. This paper is organized as follows: In Sec. II, we outline the formalism for computing decay widths within the QCD sum rule approach. Sec. III presents the numerical results for the $J^{P}=1/2^{-}$ pentaquark’s decay properties. Sec. IV discusses its production mechanism via $\Xi_{bc}^{+}$ decays and estimates detectable branching ratios. A brief summary is presented in Sec. V.
@@ II Three Point QCD Sum Rule
Over past several decades, the method of QCD sum rule has been proven to be very powerful to study hadron properties Reinders:1984sr ; Shifman:1978bx ; Colangelo:2000dp ; Narison:2002woh . In this section, we shall study the three-point correlation function of several two-body strong decay processes $M\rightarrow X+Y$. For the strong decay process $M\rightarrow X+Y$, the corresponding correlator is written as
@@
where $J_{M(X,Y)}$ is the interpolating current for the initial(final) state. In this section, we shall consider the $P_{ccs}\Xi_{cc}\bar{K}$ and $P_{ccs}\Omega_{cc}\pi$ strong decay vertices with $K(\pi)$ off shell. We use the following interpolating current for $P_{ccs}^{++}$ by considering it as a $\Xi_{cc}\bar{K}$ molecule Yang:2024okq :
@@
where $\mathcal{C}$ denotes the charge conjugate operator, subscript $a\cdots d$ denotes the color index and $u,d,s,c$ denotes the up, down, strange, charm quark field, respectively. This current can couple to the $P_{ccs}^{++}$ state with $J^{P}=1/2^{-}$ via
@@
which can couple to the meson states via
@@
which can couple to the baryon states via
@@
On the OPE side, we can evaluate the correlation function with standard QCD sum rule approach. To establish a sum rule for the coupling constant, we will pick out the $1/Q^{2}$ terms around the pole $Q^{2}\sim 0$ (where $Q^{2}=-q^{2}$) with the structure $\not${p} in the OPE series and then match both sides of the sum rule. The mass of the off-shell meson ($\bar{K}$ or $\pi$) is much smaller than the masses of the baryons ($\Xi_{cc}$, $\Omega_{cc}$, $P_{ccs}$). This justifies the approximation of extracting the $1/q^{2}$ pole term directly in the QCD sum rule analysis, making an exponential extrapolation to the on-shell point unnecessary. After performing the Borel transform $P^{2}\rightarrow M_{B}^{2}$ on both phenomenological and OPE sides, we obtain the strong coupling for the $P_{ccs}\Xi_{cc}\bar{K}$ vertex
$g_{P_{ccs}\Xi_{cc}\bar{K}}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Xi_{cc}}\lambda_{K}(m_{P_{ccs}}+m_{\Xi_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Xi_{cc}}^{2}}{\mathrm{e}^{-m_{\Xi_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (13)
@@
$g_{P_{ccs}\Omega_{cc}\pi}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Omega_{cc}}\lambda_{\pi}(m_{P_{ccs}}+m_{\Omega_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Omega_{cc}}^{2}}{\mathrm{e}^{-m_{\Omega_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (14)
@@
Using the operator product expansion (OPE) method, the three-point function can also be evaluated at the quark-gluonic level as a function of various QCD parameters. To evaluate the Wilson coefficients, we adopt the heavy quark propagator in momentum space and the light quark propagator in coordinate space
@@
where $Q$ represents the heavy quark $c$ or $b$, $q$ represents the light quark $u,d,s$, the superscripts $a,b$ denote the color indices. In this work, we will evaluate Wilson coefficients of the correlation function up to dimension nine condensates at the leading order in $\alpha_{s}$. The spectral function $\rho(s)$ in Eqs. (13), (14) are given in Appendix A. We shall discuss the detail to obtain suitable parameter working regions in QCD sum rule analysis in next section.
@@ III Numerical Analysis
where $t=\mathrm{log}\frac{\mu^{2}}{\Lambda^{2}}$, $b_{0}=\frac{33-2n_{f}}{12\pi}$, $b_{1}=\frac{153-19n_{f}}{24\pi^{2}}$, $b_{2}=\frac{2857-\frac{5033}{9}n_{f}+\frac{325}{27}n_{f}^{2}}{128\pi^{3}}$, $\Lambda=$210 MeV, 292 MeV and 332 MeV for the flavors $n_{f}=$5, 4 and 3, respectively. In this work, we evolve all the input parameters to the energy scale $\mu=m_{c}$ for our sum rule analysis. The parameters adopted for the $K$ and $\pi$ mesons and double charm baryons are displayed in Tab. 1.
@@ III.1 Strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$
Figure 2: The dependence of the strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$ on the Borel mass $M_{B}^{2}$. The transfer momentum is set to be $Q^{2}=m_{\Xi_{cc}}^{2}\sim 13.1\;\mathrm{GeV}^{2}$.
In Fig. 2, we show the variation of the coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Xi_{cc}}^{2}\approx 13.1\;\text{GeV}^{2}$. This momentum point is chosen far away from the kaon pole $m_{K}^{2}$ so that the approximation of extracting the $1/Q^{2}$ term in the QCD sum rule is valid. We find that the coupling constant has a maximum value at $M_{B}^{2}\sim 1.77\;\text{GeV}^{2}$, where it shows minimal dependence on the unphysical Borel parameter. Since the kaon mass is much smaller than the masses of the baryons $\Xi_{cc}$ and $P_{ccs}$, the $Q^{2}$-dependence of the coupling constant is negligible in the range from the sum rule working point $Q^{2}=m_{\Xi_{cc}}^{2}$ to the physical on-shell point $Q^{2}=-m_{K}^{2}$. Therefore, we can directly use the coupling constant evaluated at $Q^{2}=m_{\Xi_{cc}}^{2}$ to compute the decay width without performing an extrapolation. The extracted coupling constant is:
$g_{P_{ccs}\Xi_{cc}\bar{K}}(m_{\Xi_{cc}}^{2})=-(0.45\pm 0.05)\;\text{GeV}^{-3}.$ (18)
This value is then used to compute the decay width. From the matrix element, we can obtain the decay width for $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Xi_{cc}}^{2},m_{K}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Xi_{cc}\bar{K}}(m_{\Xi_{cc}}^{2})\left((m_{P_{ccs}}+m_{\Xi_{cc}})^{2}-m_{K}^{2}\right).$ (19)
Substituting the above coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=65\pm 16\,\mathrm{MeV}.$ (20)